ganeshie8 (ganeshie8):

Suppose a less dense object is thrown into a dense pond. Would the object bounce off the water surface ?

OpenStudy (devonhoward15):

It would float... right?

OpenStudy (electricimpurity):

I don't think it would...it would sink and then come back up to the top.

OpenStudy (hyperlunatic):

i would think it would float after sinking like @ElectricImpurity said, but i don't know...

OpenStudy (l):

In an ideal situation, it'd go in, come back up and bounce out of the surface with the same speed with which it got in. Unfortunately, factors like viscosity don't allow that to happen.

OpenStudy (electricimpurity):

Wouldn't it (at least) slightly depend on the shape of the object?

OpenStudy (l):

Assuming that only buoyancy is a factor, no.

OpenStudy (ashblood0323):

I agree. when a less dense object hits the dense pond it will sink and come and sit on the top of the water.

ganeshie8 (ganeshie8):

Viscosity only dampens the amplitude I guess. If the system is not overdamped, the object should make noticeable oscillations before dying?

OpenStudy (sshayer):

a part of the object will immerse in the water and rest will float. such that mass of amount of water displaced=mass of the object.

ganeshie8 (ganeshie8):

Viscosity only dampens the amplitude I guess. If the system is not overdamped, the object should make noticeable oscillations, bouncing off the water surface and sinking down periodically before dying ?

OpenStudy (l):

Let's say that it's thrown into the water with a velocity \(v\) vertically. Also suppose that it accelerates up with an acceleration \(a\) due to the buoyant force. The depth that it'd travel up to is given by\[0^2 = v^2 - 2ad \Rightarrow d = \frac{v^2}{2a}\]Now it starts at that depth and travels up with an acceleration of \(a\). At the top, the velocity would be given by\[v_{top}^2 = 2da = 2a\cdot \frac{v^2}{2a} = v^2\]\[v_{top} = v\]So the velocity at the top would be the velocity with which the stone is thrown in. And then we all know that the stone would simply bounce off the surface of water with the same velocity with which it was thrown. So no dampening in the amplitude of the oscillation in very, very ideal conditions.

ganeshie8 (ganeshie8):

How do we know that viscosity eats away all the ke of the object?

OpenStudy (l):

That's simply the work-energy theorem in action.

ganeshie8 (ganeshie8):

We have both posted the replies at the same time. My earlier reply is related to my previous reply to that...

ganeshie8 (ganeshie8):

For sure ideal conditions do no good here. We cannot ignore viscosity as we don't see objects bouncing off water surface usually...

OpenStudy (l):

Oh, btw, are you considering the presence of viscosity? There's an added Stokes Theorem for that.

ganeshie8 (ganeshie8):

I'm gonna look it up... Have you ever seen any object bouncing off water surface ?

ganeshie8 (ganeshie8):

Looks the object will continue sinking down till all its ke is eaten by viscosity. If the object's density is same as the water's, it stays there. Otherwise it sinks all the way down or rises to the surface due to gravity/buoyancy.

OpenStudy (mathmate):

Sorry to barge in. Both cases are possible, depending on the difference in density, viscosity (air and liquid), shape factor of the object, etc. Example that sink and come back up: a swimmer who dives in a pool. Example that does not sink at all: throw a styrofoam ball from 10 ft above water.

ganeshie8 (ganeshie8):

Very interesting examples I think the styrofoam will shoot out of the water surface if we release it from 100 feet down the water surface. Even though buoyancy continually accelerates it up, it might reach the terminal velocity due to viscosity before reaching the water surface.

ganeshie8 (ganeshie8):

Oh I forgot gravity

OpenStudy (mathmate):

True, never thought of terminal velocity going UP!

ganeshie8 (ganeshie8):

Nope. Buoyancy accounts for gravity. So above reply is perfectly fine I guess

ganeshie8 (ganeshie8):

My nope is not for your reply :)

OpenStudy (mathmate):

One more variable: angle of approach. Think of the beach where you'd throw a flat piece of stone almost tangent to the surface of the water. It doesn't sing (right away).

ganeshie8 (ganeshie8):

Honestly I had never thought about the physics of stone skipping on water surface!

OpenStudy (nincompoop):

I have seen objects bounce off water.

ganeshie8 (ganeshie8):

skipping stones @nincompoop ?

OpenStudy (nincompoop):

no

OpenStudy (nincompoop):

surface tension (upward force created by the film of water) will allow objects to bounce off in some permissible conditions

OpenStudy (nincompoop):



OpenStudy (bobo-i-bo):

https://www.youtube.com/watch?v=k4767cZzS28

OpenStudy (bobo-i-bo):

Looks like a lot of dynamics and properties of things are involved O.o

ganeshie8 (ganeshie8):

Holy balls!

OpenStudy (nincompoop):